During 0.26 s, a wheel rotates through an angle of 2.44 rad as a point on the periphery of the wheel moves with a constant speed of 2.27 m/s. What is the radius of the wheel?
Answer #1

We know that
v = wr

Thus r = v/w

Here w = 2.44/0.26 = 9.38 rad/sec

 

Thus r = 2.27/9.38 = 0.242 m